# Let’s Verify Math Questions Step by Step

### v1

@@ 3.2 Question Correctness Verification
Figure 2: An example of an incorrect atomic condition: a negative area contradicts fundamental mathematical definitions.
Step4: Contradictions among Conditions.
To assess the internal consistency of a math question $q_{i}$, we examine all subsets of its atomic conditions $\mathcal{P}(q)$ (with subset size $|S|$, $2\leq|S|\leq N$), ensuring no logical contradictions arise when conditions are considered jointly. At this stage, each condition is individually valid, and pairwise consistency has been verified; this step ensures global coherence across the full set. Formally, we require:
[…]
Step5: Condition Completeness Evaluation.
After verifying the absence of contradictions among atomic conditions, we assume the condition space $\mathcal{P}(q_{i})$ is internally consistent. Under this assumption, if a target query $G_{i}\in\mathcal{G}(q_{i})$ remains unsolvable, it indicates that essential information is missing. We formalize this by checking whether each goal is logically derivable from the available conditions:
$\text{Complete}(q_{i})=1\iff\forall G_{i}\in\mathcal{G}(q_{i}),\quad\text{Derivable}(G_{i}\mid\mathcal{P}(q_{i}))=1$
Redundant but non-conflicting conditions are permitted, as they do not compromise derivability. This criterion ensures that the classification of a question as under-specified stems solely from genuine informational gaps, not from internal inconsistency.
Combining all verification steps, the full verification function is given by:
$\mathcal{V}(q_{i})=\text{InstValid}(q_{i})\land\text{Clean}(q_{i})\land\text{Valid}(q_{i})\land\text{ContradictionFree}(q_{i})\land\text{Complete}(q_{i})$
3.3 Multi-Model Voting Strategy
To enhance the robustness of condition verification, we propose a multi-model voting strategy that aggregates predictions from multiple independently trained models. Instead of relying on a single model’s output, this strategy leverages majority voting to reach a collective decision. Formally, we consider configurations denoted as $(n,k)$, where $n$ is the total number of models in the ensemble, and $k$ is the minimum number of models that must agree (i.e., output a positive decision) for the final prediction to be accepted.
$V^{\text{final}}(q_{i})=\begin{cases}1,&\text{if }\sum\limits_{j=1}^{n}V_{j}(q_{i})\geq k\\ 0,&\text{otherwise}\end{cases}$ (1)
@@ 3.3 Multi-Model Voting Strategy
Table 1: MathQ-Verify and baseline performance on MathClean-GSM8K and MathClean-MATH.
Model Method MathClean-GSM8K MathClean-MATH
@@
4 ValiMath: A Step-Tagged Extension of MathClean
While MathClean[19] provides a benchmark for validating math problem correctness, it primarily contains simpler, synthetic questions and lacks fine-grained annotations. To introduce a more diverse range of difficulties for comprehensive evaluation, we present ValiMath, a benchmark of 2,147 questions (1,299 correct, 848 incorrect) covering five diverse error types as shown in Table 5.
Our dataset encompasses a wide range of difficulty levels and diverse mathematical domain labels. The validity of the data is ensured through a rigorous filtering and validation process. We have reorganized and developed a new secondary classification system based on the MSC classification standard [3], which has been systematically applied to our data. The dataset is constructed through a meticulous multi-stage pipeline, involving mathematics domain experts to guarantee high-quality annotations. This comprehensive process ensures both the correctness of the questions and the accuracy of the labels. More details and the complete data construction process, along with data analysis results, are provided in Appendix A.
5 Experiments
In this section, we conduct a series of rigorous experiments to systematically evaluate the effectiveness and impact of our proposed pipeline. Specifically, we aim to address the following research questions:
Q1: How does MathQ-Verify perform on established benchmarks, and is the proposed pipeline adaptable across different model architectures?
@@ 5 Experiments
5.1 Experimental Setup
Models.
We evaluate a total of 14 large language models in this study, encompassing both reasoning-oriented and non-reasoning models. The reasoning-oriented models, such as DeepSeek-R1-671B, Gemini-2.5-Pro, QwQ-32B and GPT-o4-mini, are explicitly optimized for handling complex tasks that require multi-step inferential reasoning. In contrast, the non-reasoning models, including Qwen2.5-7B, Qwen2.5-Math-7B, Qwen2.5-72B, Qwen2.5-Math-72B, Llama-3.1-8B, S1.1-32B, Llama-3.1-70B, and GPT-4o, are primarily designed for general-purpose language understanding or specialized mathematical reasoning without explicit support for extended reasoning chains.
Dataset.
We utilize the synthesized and annotated versions of GSM8K and MATH from the MathClean benchmark as our primary evaluation datasets. Additionally, we incorporate the ValiMath dataset, as described in Section 4, to comprehensively assess the performance of our approach.
Table 2: Comparison of our MathQ-Verify and baseline performance on ValiMath.
@@ 5.1 Experimental Setup
We adopt standard evaluation metrics to assess model performance, including accuracy, precision, recall, F1 score, the number of invalid outputs, and step-wise accuracy. Accuracy reflects the overall correctness, while precision and recall measure the proportion of correct predictions among all valid outputs and among all answerable questions, respectively. The F1 score provides a balanced summary of these two metrics. Invalid outputs refer to the number of samples from which no valid answer can be extracted using either instruction-based or rule-based decoding; final judgments on invalidity are made through semantic evaluation using GPT-4o. Step-wise accuracy measures the correctness of each individual step executed in the MathQ-Verify framework. S1&2 to S5 represent the accuracy of predictions for Question correctness at each steps in the ValiMath. Notably, under comparable recall, our method consistently achieves higher precision, indicating stronger robustness in identifying valid questions and resisting spurious outputs.
Baseline.
@@
For the analysis of invalid samples on the MathClean-Benchmark, we adhere to the evaluation methodology established in the original work. For the ValiMath dataset, we construct a baseline method that assesses the correctness of each input question directly, without employing the decomposition or multi-step verification procedures integral to the MathQ-Verify framework. This baseline provides a point of comparison to evaluate the efficacy of our proposed pipeline.
Settings.
@@
All open-source models are deployed using the vLLM framework, while closed-source models are accessed via their respective official APIs. For open-source models, we adopt consistent settings: temperature=0.7, top_p=0.8, top_k=20, and max_tokens=16384. For math-specialized models (e.g., Qwen2.5-Math-7B), we set max_tokens=4096 to optimize performance. All experiments are conducted on 8 x NVIDIA H200 GPUs.
5.2 Results and Analysis
@@
To address Q1, which investigates the effectiveness of the proposed MathQ-Verify framework, we conduct comprehensive evaluations on three benchmark datasets: MathClean-GSM8K, MathClean-MATH, and ValiMath. The primary objective is to quantify the extent to which MathQ-Verify improves prediction accuracy and to assess its applicability across diverse model architectures. As presented in Table 1, our method yields substantial improvements in both accuracy and the reduction of invalid outputs across a wide range of models. In particular, models such as Qwen2.5-7B and LLaMA-3.1-8B exhibit notable gains in precision and F1 scores. However, certain specialized models such as Qwen2.5-Math-7B demonstrate limitations in instruction-following capabilities, resulting in increased invalid outputs. These observations suggest that models with sufficient instruction-following capabilities generally exhibit performance improvements when augmented with MathQ-Verify.
To further evaluate the robustness and generalizability of MathQ-Verify, we extend our analysis beyond the MathClean benchmarks to the ValiMath dataset. This investigation aims to determine whether the observed performance gains persist under distributional shifts and dataset variations. As illustrated in Table 2, MathQ-Verify consistently enhances performance across datasets. For example, GPT-4o and S1.1-32B achieve an improvement of 5.78 and 6.02 percentage points in F1 score, respectively, alongside increases in precision. These findings affirm the framework’s ability to generalize across diverse data distributions, effectively improving both the reliability and precision of model outputs while mitigating spurious generations.
Furthermore, we conducted a stepwise evaluation from S1 to S5 to assess the contribution of each step to overall performance. This evaluation aimed to understand how the various stages of the pipeline interact with different model and how they influence performance. The results indicate that both non-reasoning and reasoning models exhibit improvements in the early stages (S1&2 and S3), with notable gains observed in models such as Qwen2.5-7B. However, reasoning models, such as QwQ-32B, showed a decline in accuracy during the later stages (S4 and S5). This step-wise analysis provides insights into how MathQ-Verify interacts with pipeline stages and reveals that its effectiveness is influenced by the model’s reasoning capabilities and task complexity.
Figure 3: Distribution heatmaps of correct questions across difficulty levels and primary mathematical categories. The left heatmap shows human-annotated correct question distribution, while the right heatmap shows the distribution of questions validated as correct by MathQ-Verify after filtering.
5.3 Multi-Model Voting Ensemble Trade-Offs.
In this subsection, we aim to address Q2, whether majority voting can further enhance precision. The numbers correspond to the models as indicated in Table 2.
Considering the trade-offs between precision and recall, and the objective of retaining a broad and diverse set of high-quality samples, the $(2,2)$ configuration emerges as the optimal choice. It provides an effective balance, enhancing mathematical problem verification while ensuring that recall remains acceptable. Different configurations support varying precision and cost needs. This approach highlights the flexibility of $(n,k)$ voting, enabling us to prioritize precision without sacrificing recall, and ultimately improving the overall performance of the pipeline.
Table 3: Different model ensembles for $(n,k)$ configurations.
@@
5.4 Consistency of Data Distribution
To answer Q3, which examines whether MathQ-Verify introduces significant distributional bias, we compare the joint category–difficulty distributions of the original human-annotated correct questions with those filtered by MathQ-Verify, as shown in Figure 3. The heatmaps reveal that the filtered distribution closely matches the original, with only minor shifts observed in certain categories and difficulty levels. Low-density areas, such as extreme difficulty levels and peripheral categories, remain consistent in both distributions. These results indicate that MathQ-Verify preserves the original category–difficulty structure, maintaining consistency with the human-annotated distribution.
5.5 Ablation Study
@@
Table 4: Data distribution of ValiMath.
Category Count %
Instruction Error 143 6.66
Nonsemantic Check 119 5.54
Minimal Domain Check 171 7.96
Contradiction Check 300 13.97
Completeness Check 115 5.36
Correct Instances 1299 60.50
Table 5: Ablation study.
@@
6 Conclusion
In this paper, we introduce a high-precision filtering pipeline for mathematical questions in LLM training data, in contrast to prior work that predominantly filters based on answer correctness. First, we perform format-level, linguistic filtering to isolate well-formed mathematical questions. Next, we formalize each problem by defining atomic conditions and inter-condition consistency checks such as domain conformity and contradiction detection allowing us to decompose each question into verifiable units. Finally, a goal-driven completeness assessment rigorously identifies any missing conditions. To facilitate fine-grained, step-wise validation, we also construct the MVQ-2K benchmark, comprising 2147 human-verified examples. Experiments on different LLMs show that a single MathQ-Verify model achieves state-of-the-art performance on MathClean and MVQ-2K, and that merely a majority-voting method raises precision up to 92%. Our pipeline thus provides a robust, high-accuracy solution for constructing quality mathematical QA datasets.
Appendix A The Construction Process of ValiMath
Figure 4: Overview of the ValiMath Construction Pipeline
Synthesis.
For synthetic data in the mathematics domain, especially for the generation of questions, there is a high demand for diversity. Additionally, the synthetic data itself tends to align more closely with the distribution of model training data. To construct a diverse set of questions, while also introducing more challenging problems compared to MathClean [19], we thoroughly researched existing prompts used for generating synthetic questions. Based on our findings, we carefully designed a prompt that ensures the diversity of the generated questions. To ensure the generated mathematical problems possess a certain level of difficulty and challenge, we ultimately used NuminaMath [18] as our source data for question synthesis.
Furthermore, to capture a broad range of potential errors that the model might generate, we employed models of different sizes for question synthesis. Stronger reasoning models tend to generate problems with a lower error rate during normal question synthesis, while smaller models are more likely to produce erroneous questions. Therefore, we used a variety of models, including: Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-72B-Instruct, GPT-4o, and GPT-o3mini, ensuring that each model generated a similar number of questions. In total, we synthesized 10,000 mathematical questions.
Filtering.
To ensure the uniqueness and determinacy of the final answers, we applied a GPT-4o-based prompt filtering process to exclude all multiple-choice questions. It is important to note that, since our objective is to construct a dataset containing both correct and incorrect problems, no additional filtering was performed beyond this step. This preserves the natural distribution of model-generated outputs.
Data Selection.
To reduce annotation costs in advance, we observed during preliminary experiments that general-purpose models with strong reasoning ability—such as GPT-4o tend to generate more correct than incorrect mathematical problems in our raw data (with fewer than 30% incorrect samples). To address this, we first performed a coarse-grained filtering step to retain a higher proportion of incorrect problems for annotation.
Specifically, we employed Qwen2.5-72B-Instruct to directly assess the correctness of each generated problem using the baseline judgment method described in the main paper. Based on these results, we curated a subset of approximately 4,000 mathematical problems with a target ratio of correct to incorrect samples set at 1:3. This step effectively reduced downstream annotation workload while preserving as many erroneous problems as possible, thereby enhancing the effectiveness of subsequent data verification.
Expert Annotation.
After obtaining the initial set of 4,000 mathematical problems, we invited domain experts in mathematics to perform detailed annotations. The goal was to ensure the accuracy of the binary correctness labels ("correct" or "incorrect") and, for incorrect problems, to further annotate the specific steps where errors occurred, following the five-stage framework defined in MathQ-Verify. The annotation process proceeded as follows:
  1. 1.
    Annotation Guidelines. Experts were provided with comprehensive annotation instructions, including precise definitions of each step in the MathQ-Verify process as outlined in the main paper. Each problem was evaluated across the five reasoning stages, and labeled accordingly as either “correct” or “incorrect.” For incorrect problems, annotators further specified the earliest step at which the error occurred.
  2. 2.
    Independent Dual Annotation. Each problem was independently annotated by at least two domain experts to mitigate individual biases. Discrepancies between annotations were resolved through a final joint review process, during which experts reached a consensus based on predefined labeling criteria.
  3. 3.
  4. 4.
    Subjectivity Analysis. To further assess and control for annotator subjectivity, we randomly sampled a subset of annotated problems for inter-annotator agreement analysis. We applied standard consistency metrics to quantify labeling stability across experts and identify any systemic annotation biases.
  5. 5.
    Quality Assurance. Through the above multi-step annotation workflow, we ensured the reliability, consistency, and accuracy of the labeled dataset, providing a high-quality foundation for downstream evaluation and model training.
Appendix B Additional Implement Details
B.1 Question Labels

### v2

@@ 4.1. Problem Formulation
Figure 2. An example of an incorrect atomic condition: a negative area contradicts mathematical definitions.
4.2. Question Correctness Verification
We verify the correctness of a mathematical question $q$ using a five-stage pipeline, MathQ-Verify, which decomposes $q$ into formal components and checks them for instruction integrity, linguistic soundness, atomic validity, global consistency, and goal completeness. This subsection provides (i) precise notation, (ii) five predicates that formalize each stage, and (iii) a single overall decision rule $V(q)\in\{0,1\}$, Formal algorithm is provided in Appendix C.
Step 1: Contaminated Instruction Detection. We first ensure that $q$ is a genuine math question and that its phrasing does not bias, leak, or otherwise contaminate the problem-solving objective. Typical violations include meta-instructions (e.g., “please rewrite” / “please rephrase”) and explicit answers embedded in the prompt. Formally, we define $\mathrm{InstValid}(q)=True$ if $q$ does not have any violation instructions above.
Step 2: Linguistic Error Detection. A question $q$ is considered clean if it has no semantic imperfections, including spelling mistakes, grammar errors, and Latex anomalies, that hinder readability or parsing while leaving the mathematical content intact. Then we define $\mathrm{Clean}(q)=True$ if there is no semantic imperfections mentioned above.
Step 3: Atomic Condition Error Detection. We decompose a ”clean” (defined in Step 2) question into several minimal conditions, called atomic conditions, and do some checks below directly using prompting LLM. The prompt is provided in the repository. Each atomic condition $P_{j}\in\mathcal{P}(q)$ must conform to the basic rules of the relevant domain (algebra, geometry, analysis, etc.). We check atomic validity at the statement level (e.g., rejecting definitions that contradict standard semantics, such as treating a discrete variable as continuous without qualification). If all the conditions in $q$ obeys all these mathematical rules above, then $\mathrm{AtomValidAll}(q)=True$.
Step 4: Global Consistency Checking. Beyond the validity of individual atoms, the full condition set $\mathcal{P}(q)$ must be jointly satisfiable. Instead of quantifying over all nontrivial subsets (which is cumbersome and unnecessary), we require a single global condition $\mathrm{Sat}(\mathcal{P}(q))$. By prompting LLM (provided in repository), if the model outputs “no contradictions” at the level of the entire premise set then we define $\mathrm{Consistent}(q)=True$.
Step 5: Condition Completeness Evaluation. After obtaining the consistent $\mathcal{P}(q)$, we check the completeness by LLM (provided in repository) that every target in $\mathcal{G}(q)$ should be derivable from the available conditions, then $\mathrm{Complete}(q)=True$. If some $G\in\mathcal{G}(q)$ is not entailed (cannot derive answer from premises), then the question is under-specified (missing information).
Therefore, the overall decision for q is determined as follows:
(1) $V(q)$ = $=\mathbf{1}\Big[\mathrm{InstValid}(q)\ \wedge\ \mathrm{Clean}(q)\ \wedge\ \mathrm{AtomValidAll}(q)\$
$\quad\quad\quad\wedge\ \mathrm{Consistent}(q)\ \wedge\ \mathrm{Complete}(q)\Big]$
The following are some illustrative examples.
  1. (1)
    Contradiction (fails Step 4). Let $\mathcal{P}(q)=\{x\in\mathbb{Z},\ x\ \text{is even},$ x=2k+1}\ x=2k+1\} with $k\in\mathbb{Z}$. The set is not satisfiable, so $\mathrm{Consistent}(q)$ fails, hence $V(q)=0$ by (1).
  2. (2)
    Under-specification (fails Step 5). If $\mathcal{P}(q)=\{\text{``$n$ is a}\text{polygon''}\}$ and $\mathcal{G}(q)=\{\text{``compute the sum of interior angles''}\}$, but the $\mathcal{P}(q)$ does not have the numeric value of polygon sides. Thus $\mathrm{Complete}(q)$ fails and $V(q)=0$.
  3. (3)
    Valid (passes all steps). If $\mathcal{P}(q)=\{x\in\mathbb{Z},\ x\ \text{is even},\ 2x+3=19\}$ and $\mathcal{G}(q)=\{\text{solve for }x\}$, then the atoms are valid and jointly satisfiable; moreover, $x=8$ is derivable from $\mathcal{P}(q)$. Hence $V(q)=1$.
Remarks on implementation. The five predicates abstract away from any particular detector or solver; concrete implementations (e.g., rule-based checks, symbolic solvers, or verifiers) can be described in the implementation section. Keeping the formal layer model-agnostic improves clarity and reproducibility, while (1) provides a single, auditable criterion for acceptance.
4.3. Multi-Expert Voting Strategy
@@
Table 2. MathQ-Verify and baseline performance on MathClean-Simple and MathClean-Challenging.
Inv. denotes answers requiring GPT-4o-based semantic parsing; green highlights indicate the best scores, same for later tables.
Model Method MathClean‐Simple MathClean‐Challenging
@@
To enhance the robustness of condition verification, we propose a multi-expert voting strategy that aggregates predictions from multiple independently models. Instead of relying on a single model’s output, this strategy leverages majority voting to reach a collective decision. Formally, we consider configurations denoted as $(n,k)$, where $n$ is the total number of models in voting, and $k$ is the minimum number of models that must agree (i.e., output a positive decision) for the final prediction to be accepted.
(2) $V^{\text{final}}(q_{i})=\begin{cases}1,&\text{if }\sum\limits_{j=1}^{n}V_{j}(q_{i})\geq k\\ 0,&\text{otherwise}\end{cases}$
5. Experiments
In this section, we conduct a series of rigorous experiments to systematically evaluate the effectiveness and impact of our pipeline. Specifically, we aim to address the following research questions:
Q1: How does MathQ-Verify perform on established benchmarks, and is the pipeline adaptable across different model architectures?
@@ 5. Experiments
5.1. Experimental Setup
Models. We evaluate a total of 14 large language models in this study, encompassing both reasoning-oriented and non-reasoning models. All models are Instruct version if only labeled with parameter size. The reasoning-oriented models, such as DeepSeek-R1-671B, Gemini-2.5-Pro, QwQ-32B and GPT-o4-mini, are explicitly optimized for handling complex tasks that require multi-step inferential reasoning. In contrast, the non-reasoning models, including Qwen2.5-7B, Qwen2.5-Math-7B, Qwen2.5-72B, Qwen2.5-Math-72B, Llama-3.1-8B, S1.1-32B, Llama-3.1-70B, and GPT-4o, are primarily designed for general-purpose language understanding or specialized mathematical reasoning without explicit support for extended reasoning chains.
Dataset. We utilize the synthesized and annotated versions of mathematical data from the MathClean benchmark as our primary evaluation datasets. Additionally, we incorporate the ValiMath dataset, as described in Section 3, to comprehensively assess the performance of our approach.
Table 3. Comparison of our MathQ-Verify and baseline performance on ValiMath.
Skk denotes evaluating the accuracy as the same of $k$-th step to the entire dataset directly, rather than checking step by step.
@@ 5.1. Experimental Setup
Evaluation Metrics. We adopt standard evaluation metrics to assess model performance, including accuracy, precision, recall, F1 score, the number of invalid outputs, and step-wise accuracy. Accuracy reflects the overall correctness, while precision and recall measure the proportion of correct predictions among all valid outputs and among all answerable questions, respectively. The F1 score provides a balanced summary of these two metrics. Invalid outputs refer to the number of samples from which no valid answer can be extracted using either instruction-based or rule-based decoding; final judgments on invalidity are made through semantic evaluation using GPT-4o. This design accommodates models specialized in specific tasks, aiming to explore whether they truly possess the ability to identify erroneous questions. Step-wise accuracy measures the correctness of each individual step executed in the MathQ-Verify framework. S1&2 to S5 represent the accuracy of predictions for Question correctness at each steps in the ValiMath. Notably, under comparable recall, our method consistently achieves higher precision, indicating stronger robustness in identifying valid questions and resisting spurious outputs.
@@
Baseline. For the analysis of invalid samples on the MathClean-Benchmark, we adhere to the evaluation methodology established in the original work. For the ValiMath dataset, we construct a baseline method that assesses the correctness of each input question directly, without employing the decomposition or multi-step verification procedures integral to the MathQ-Verify framework. This baseline provides a point of comparison to evaluate the efficacy of our proposed pipeline.
@@
Settings. All open-source models are deployed using the vLLM framework, while closed-source models are accessed via their respective official APIs. For open-source models, we adopt consistent settings: temperature=0.7, top_p=0.8, max_tokens=16384, top_k=20. For math-specialized models (e.g., Qwen2.5-Math-7B), we set max_tokens=4096 to optimize performance. All experiments are conducted on 8 x NVIDIA H200 GPUs.
5.2. Results and Analysis
@@
To address Q1, which investigates the effectiveness of MathQ-Verify, we conduct comprehensive evaluations on three benchmark datasets: MathClean-Simple, MathClean-Challenging, and ValiMath. The primary objective is to quantify the extent to which MathQ-Verify improves prediction accuracy and to assess its applicability across diverse model architectures. As presented in Table 2, our method yields substantial improvements in both accuracy and the reduction of invalid outputs across a wide range of models. In particular, models such as Qwen2.5-7B and LLaMA-3.1-8B exhibit notable gains in precision and F1 scores, which reflects the likelihood that predicted samples contain truly correct data. However, certain specialized models such as Qwen2.5-Math-7B demonstrate limitations in instruction-following capabilities, resulting in increased invalid outputs. These observations suggest that models with sufficient instruction-following capabilities generally exhibit performance improvements when augmented with MathQ-Verify.
Table 4. Different model ensembles for $(n,k)$ configurations.
TP indicates correct positive predictions, while FP indicates incorrect positive predictions.
@@
To further evaluate the robustness and generalizability of MathQ-Verify, we extend our analysis beyond the MathClean benchmarks to the ValiMath dataset. This investigation aims to determine whether the observed performance gains persist under distributional shifts and dataset variations. As illustrated in Table 3, MathQ-Verify consistently enhances performance across datasets. For example, GPT-4o and S1.1-32B achieve an improvement of 5.78 and 6.02 percentage points in F1 score, respectively, alongside increases in precision. These findings affirm the framework’s ability to generalize across diverse data distributions, effectively improving the reliability and precision of model outputs while mitigating spurious generations.
Furthermore, we conducted a stepwise evaluation from S1 to S5 to assess the contribution of each step to overall performance. For example, S3 means we directly evaluate the accuracy of the ”Atomic Condition Error” on the entire dataset to assess the ability of models capturing this type of error, rather than evaluate a sample from S1 to S3 and calculate a multi-step accuracy. This evaluation aimed to understand how the various stages of the pipeline interact with different model and how they influence performance. The results indicate that both non-reasoning and reasoning models exhibit improvements in the early stages (S1&2 and S3), with notable gains observed in models such as Qwen2.5-7B. However, reasoning models, such as QwQ-32B, showed a decline in accuracy during the later stages (S4 and S5). Note that the decrease here refers to performance relative to the baseline in detecting this error type; it shows that some models perform worse with more stepwise hints. S5 reflects only the accuracy of directly detecting this error, unrelated to full question verification accuracy. This step-wise analysis provides insights into how the every step in MathQ-Verify performs capturing different kinds of errors.
Figure 3. Distribution heatmaps of correct questions across difficulty levels and primary mathematical categories. The left heatmap shows human-annotated correct question distribution, while the right heatmap shows the distribution of questions validated as correct by MathQ-Verify after filtering.
5.3. Multi-Model Voting Ensemble Trade-Offs.
In this subsection we aim to address Q2, The numbers correspond to the models as indicated in Table 3.
More voters means the increase of cost. Considering the trade-offs between precision and recall, and the objective of retaining a broad and diverse set of high-quality samples, the $(2,2)$ configuration emerges as the optimal choice. It provides an effective balance, enhancing question verification while ensuring that recall remains acceptable. Different configurations support varying precision and cost needs. This approach highlights the flexibility of $(n,k)$ voting, enabling us to prioritize precision without sacrificing recall, and ultimately improving the overall performance of the pipeline.
5.4. Consistency of Data Distribution
To answer Q3, which examines whether MathQ-Verify introduces significant distributional bias, we compare the joint category and difficulty distributions of the original human-annotated correct questions with those filtered by MathQ-Verify, as shown in Figure 3. The heatmaps reveal that the filtered distribution closely matches the original, with only minor shifts observed in certain categories and difficulty levels. Low-density areas, such as extreme difficulty levels and peripheral categories, remain consistent in both distributions. These results indicate that MathQ-Verify preserves the original category–difficulty structure, maintaining consistency with the human-annotated distribution.
5.5. Ablation Study
@@
Table 5. Ablation study.
@@
6. Conclusion
In this paper, we introduce a five steps math question-filtering pipeline in LLM training data, in contrast to prior work that predominantly filters based on answer correctness. To facilitate fine‐grained, step‐wise validation, we also construct the ValiMath benchmark, comprising 2,147 human‐verified examples. It is the first publicly released benchmark that performs step‑by‑step error checking on mathematical questions; every instance is annotated by multiple domain experts, providing new insights for the open‑source community on how to curate higher‑quality datasets. Experiments on different LLMs show that a single MathQ-Verify model achieves state‐of‐the‐art performance on different benchmarks. Both our five‑stage filtering pipeline and the expert‑annotated ValiMath benchmark offer the community a turnkey toolset, a scalable way to cleanse existing corpora and a gold‑standard testbed for validity‑aware model development.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (92470121, 62402016), National Key R&D Program of China (2024YFA1014003), Humanities and Social Sciences Research Planning Fund Project of the Ministry of Education: ”Research on Metacognitive Diagnosis Theory and Technology Driven by Multimodal Learning Data” (23YJA880091), and High-performance Computing Platform of Peking University.
Appendix A Additional Implement Details
A.1. Question Labels